Fictitious element X has an a average atomic mass of 254.9 and only 2 isotopes, One of its isotopes has an abundance of 72.00 and a mass of 250.9 amu. What is the atomic mass of the second isotope?
1
Expert's answer
2020-12-11T02:50:45-0500

Solution:

The average atomic mass of an element is the sum of the masses of its isotopes, each multiplied by its natural abundance.

Average atomic mass = f1M1 + f2M2 +… + fnMn where f is the fraction representing the natural abundance of the isotope and M is the mass number (weight) of the isotope.

In our case: Average atomic mass = f1M1 + f2M2


Average atomic mass = 254.9 amu

f1 = 72.00% or 0.72

M1 = 250.9 amu

f2 = (1 - f1) = (1 - 0.72) = 0.28

M2 = unknowm


Hence,

Average atomic mass of X = f1M1 + f2M2

254.9 amu = (0.72 × 250.9 amu) + (0.28 × M2)

254.9 amu = 180.648 amu + 0.28 × M2

M2 = (254.9 amu - 180.648 amu) / 0.28 = 265.186 amu = 265.2 amu

M2 = 265.2 amu


Answer: The atomic mass of the second isotope is 265.2 amu.